[Hallc_running] [New Logentry] RC Daily Update

pooser at jlab.org pooser at jlab.org
Tue Nov 20 09:55:02 EST 2018

Logentry Text:
RC report:
We began yesterday taking data on kinematic group 13 and had stable beam for the remainder of the day and night (https://logbooks.jlab.org/entry/3628874).  Mike T. noticed that the archived Hall-C beam momentum measurement was saturating (https://logbooks.jlab.org/entry/3628448, https://logbooks.jlab.org/entry/3628464, https://logbooks.jlab.org/entry/3628489) as a result of the FFB being variously in 'SEE Normal Ops' and 'Feedback on' mode and the 'gain mode' being configured to manual.  This has been the case since 08:55 on 11/7/18 however it was remedied on 11/18/18.  Since the momentum measurement was present for currents < 40 uA and we have been consistently interspersing various targets and currents, we can simply extrapolate from the lower current running without error.

At 10:53 we HMS Q3 and dipole initiated a fast dump (https://logbooks.jlab.org/entry/3628494, https://logbooks.jlab.org/entry/3628501) as a result of a disruption to the current lead mass flows.  We recovered within an hour and continued on with group 13.  At ~16:30 we completed group 13 and changed configuration for group 16 and began collecting data 17:00.  

Data collection on group 16 remained steady throughout swing and owl shifts with no downs > 30 minutes.  At ~09:30 we completed group 16 and are currently switching polarity for group 15.  We are utilizing this polarity change in order to facilitate Jay’s beam steering tests in which we will ramp down SHMS HB and Q2 first, followed by the rest of the magnets.  We will then follow the same order in ramping them back up.  Once all magnets are ready, we will then commence with group 15.

Notices in effect for shift crews:
Shift leaders, please fill out the check list once per shift!
It is no longer required to fill out the standard.kinematics file online, it is now done automagically
Split the large desired charge between multiple runs (~1 hour per run)
Listed currents are estimates. For runs at less than 70uA, increase current until SHMS singles rates are between 450k-500k.
How-to power cycle PLC I/O chassis: https://logbooks.jlab.org/entry/3625862, https://logbooks.jlab.org/entry/3625900

Planned accelerator configuration changes:
Sunday, Nov 25(?): Hall A goes offline and MCC wants to conduct high current beam study of energy spread at 40 uA.  It is unclear that we will be effected.
Monday, Nov 26: Energy change to 1.86 GeV/pass, we will be moved over to 500 MHz laser.

Planned accesses to the hall:
Monday, Nov 26: aerogel change and HGC work needs to be coordinated with Walter K., we must be in restricted access for this work

Special requests:
Harp scan at 3H07A and 3H07B with raster off to address possible issues with harp (https://logbooks.jlab.org/entry/3624463)
Vijay Kumar (Regina group) has requested a walkthrough during the next access


This is a plain text email for clients that cannot display HTML.  The full logentry can be found online at https://logbooks.jlab.org/entry/3628894
-------------- next part --------------
An HTML attachment was scrubbed...
URL: <https://mailman.jlab.org/pipermail/hallc_running/attachments/20181120/930963d5/attachment.html>

More information about the Hallc_running mailing list